The integral int_{a}^{b} \frac{dx}{(x - a)p} diverges if . Select one: a. 2p \le 1 b. p\ge 1 c. p> 1 d. p < 1
Answer #1

Consider the integral \int_{a}^{b}\frac{dx}{(x-a)^p} .

We can evaluate the integral as

\int_{a}^{b}\frac{dx}{(x-a)^p}=\left [(x-a)^{1-p} \right ]_{a}^{b}

The integral diverges if

1-p\geqslant 0\\ {\color{Blue} p\geqslant 1}

Since, the choices are not visible, please choose p\geqslant 1 .